Circular No. 6168

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

PKS 1156+295 J. R. Webb and T. Barnello, Florida International University; I. Rob-

son, James Clerk Maxwell Telescope (JCMT); and R. C. Hartman, for the EGRET Team and multiwavelength collaboration, write: “The optically violent quasar PKS 1156+295 is currently active in the γ-ray, optical, and millimeter bands. A quick-look analysis of EGRET data for Apr. 25–29 indicated that on Apr. 27 its γ-ray emission (> 100 MeV) flared to a flux level of several × 10^{−6} photon cm^{−2} s^{−1}, one of the highest levels seen in an active galactic nucleus by EGRET. The average flux over the 4-day observation was ∼ 0.9 × 10^{−6} (same units), but it appears that most of the fluence occurred on Apr. 27. An unfiltered CCD image taken on May 2 with the SARA 0.9-m telescope at Kitt Peak indicated that PKS 1156+295 was bright in the optical region as well, with a V magnitude of 15.0 ± 0.1. Observations from the JCMT indicate that PKS 1156+295 is also bright in the millimeter region. Follow-up observations are in progress, and additional observations at all wavelengths are solicited.”

MV LYRAE S. Yu. Shugarov, Sternberg State Astronomical Institute; and V. I.

Marsakova, Odessa State University, note that MV Lyr is possibly returning from a low to an usually bright state. Their observations reveal brightness variations on Apr. 25 ranging in U between 13.90 and 14.52, in B between 15.28 and 15.88, and in V between 15.56 and 16.14. On Apr. 27, MV Lyr was much brighter (U = 13.24–13.61, B = 14.51–14.77, V = 14.68–14.75).

NO SUPERNOVA 1994ag IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports that images

taken last December with the 1-m reflector at Siding Spring confirm that the object called SN 1994ag (IAUC 6113, 6116) is a red foreground star, with no evidence of variability.

COMET 84P/GICLAS R. J. Bouma, Groningen, The Netherlands, points out that the ob-

served positions and orbital elements of comet D/1931 R1 (cf. IAUC 6161, MPC 24933, MPC 25030) are completely compatible with identity with comet 84P/Giclas.

1995 May 5 Daniel W. E. Green